OpenStudy (anonymous):

the ratio of the radii of two spheres is 6 : 5, what is the ratio of the surface areas of the two spheres?

zepdrix (zepdrix):

\[\large r_1:r_2 \qquad\rightarrow\qquad 6:5\] Surface Area of a Sphere is given by, \(\large 4\pi r^2\) So the ratio of their surface areas will be, \[\large 4\pi (r_1)^2:4\pi (r_2)^2 \qquad\rightarrow\qquad 4\pi(36):4\pi(25)\] And from there, we sould probably simplify the ratio by dividing both terms by 4pi. Yah I think we can do that... Any confusion? :o

jimthompson5910 (jim_thompson5910):

or you can use the idea that if the ratio of two sides is a/b, then the ratio of two areas is (a/b)^2 = (a^2)/(b^2)